OpenStudy (anonymous):

What is the value of the equilibrium constant for this redox reaction? 2Cr^3+(aq) + 3Sn^2+ (aq) 2Cr(s) + 3Sn^4+(aq) E = -0.89 v

OpenStudy (anonymous):

K = 5.13 x 10-93 B. K = 6.27 x 10-91 C. K = 7.92 x 10-46 D. K = 8.56 x 10-31 E. K = 9.25 x 10-16

OpenStudy (chmvijay):

E = 0.0591 /n Log Kc can u find Kc E= -0.89 and n = no of electrons involved in the reaction

OpenStudy (chmvijay):

we dont give direct answer here @david_crider we give u hint or guide u have to solve them ur self now i have give u the formula u just put them in that formula and find out

OpenStudy (anonymous):

Im in summer school for a reason haha i dont know how to apply the formula, just trying to get through this

OpenStudy (anonymous):

@chmvijay